# Subsystem Information Capacity in Random Circuits and Hamiltonian Dynamics

### v2

@@ V Discussions
Our settings are different from the conventional TMI settings where all the system qubits are entangled with corresponding reference qubits. Conversely, all qubits in $\bar{E}$ are initialized with pure states in our case. On the one hand, our settings save a great number of qubits ($L+1$ in one-to-one encoding case) and are thus much more friendly for noisy intermediate-scale quantum experiments Preskill (2018). On the other hand, the results in our settings rely on the choice of the initial states in some cases. For random quantum circuit evolution, the choice of different product initial states is irrelevant as the difference is smeared out by the first layer of random gates. Instead, the results can be different for different initial states in Hamiltonian quench dynamics. In charge conserved evolution, different choices of the initial states correspond to different charge sectors of different Hilbert space dimensions, which can result in very different behaviors Liu et al. (2024c). For different initial states within the same charge sector, the difference is less known. Therefore, it deserves further investigations into the difference in information dynamics from different initial states, particularly in the case of symmetric evolution.
@@
For U(1) symmetric dynamics, the dynamics of mutual information characterized by von Neumann entropy and Rényi entropy could be drastically different. This is because Rényi entropy growth shows subdiffusive behaviors for systems with conservation laws while von Neumann entropy always exhibits ballistic growth Rakovszky et al. (2019); Huang (2020); Zhou and Nahum (2020); Žnidarič (2020). In our work, we focus on von Neumann entropy based mutual information for Hamiltonian dynamics. The dynamical behaviors for Rényi mutual information in these systems as well as the interplay between other symmetries with information dynamics require further investigation.
@@
As previously noted, the framework for information dynamics established within this work is amenable to noisy intermediate-scale quantum Preskill (2018) technology and is thus well-suited for implementation on near-term quantum computing platforms. Specifically, our settings are more qubit efficient than measuring TMI and are free from Pauli operator average and reversal evolution than measuring OTOC. Given that $A$ and $R$ both being small regions comprising a few qubits, the von Neumann or Rényi entropy for these subsystems can be obtained either via full state tomography or random basis measurement van Enk and Beenakker (2012); Brydges et al. (2019); Huang et al. (2020). Apart from experiments on quantum hardware, an interesting theoretical future direction is to analyze the impact of different types of quantum noise on these settings utilizing statistical model mapping. This could lead to the identification of noise-resilient experimental protocols for the implementation of information dynamics. It is also worth noting that random circuit structures utilized in this work are helpful in alleviating coherent quantum noises and demonstrating quantum advantages Arute et al. (2019).
@@
This work concentrates on one-dimensional systems, and we expect that higher-dimensional systems have the potential to exhibit a broader range of behaviors in information dynamics. There are also more possible geometry configurations to arrange $E$ and $A$ regions which could lead to a richer understanding of information propagation and scrambling. Another limit is to apply information metrics developed in this work to zero-dimensional systems with all-to-all interactions such as SYK models Sachdev and Ye (1993); Maldacena and Stanford (2016); Chowdhury et al. (2022), which yield faster information scrambling and is analytically tractable in the large $N$ limit.
@@
Apart from its implications for non-equilibrium physics, SIC and a deeper understanding of information dynamics can also be helpful and insightful for designing and benchmarking quantum algorithms, particularly in the context of variational quantum algorithms Cerezo et al. (2021); Bharti et al. (2022). Promising examples along this direction include understanding the training dynamics and learning capability for variational quantum machine learning algorithms from information perspectives Shen et al. (2020); Garcia et al. (2022); Sajjan et al. (2023). Besides, it is inspiring to develop better circuit ansatz design Zhang et al. (2022); Du et al. (2022); Lu et al. (2021); Zhang et al. (2021); Wu et al. (2021), parameter initialization Park et al. (2024); Cao et al. (2024) and training strategies Liu et al. (2023d) with enhanced expressive power and alleviated barren plateau issues McClean et al. (2018) by leveraging the principles of non-equilibrium dynamics and information spreading.
@@ Acknowledgements
We thank helpful discussions with Zhou-Quan Wan and Shuai Yin.
@@ Appendix A Relation between subsystem information capacity and out-of-time correlators
The established relation between mutual information and average OTOC over local Pauli operators Hosur et al. (2016) can be similarly extended to our case:
@@
The schematic proof based on tensor network representation is shown in Fig. 29. For the tripartite mutual information setting in Ref. Hosur et al. (2016), the initial state projectors (grey triangles in Fig. 29) are replaced with direct links in pairs since the input state is a fully mixed state for each qubit. In that case, we reproduce the equivalence between Rényi-2 entanglement entropy with the conventional OTOC for local Pauli operators, unlike the introduction of projectors for the pure initial state in our settings.
@@ Appendix C Mapping between random quantum circuits and effective statistical models
The random quantum circuits under investigation are constructed from a sequence of random two-qudit unitary gates, each sampled from the Haar ensemble and arranged in a brick-wall layout, as depicted in Fig. 1(a). In this study, we adhere to the convention of defining the unit time $\Delta t$ = 1 for an even-odd layer of two-qudit gates. This convention implies a lightcone speed of $v=2$, which differs from some literature where $\Delta t=1$ is associated with a single even or odd brick-wall layer. After $T$ layers of gates, the output density matrix $\rho$ is given by
@@
where $S_{r}$ is the permutation group of degree $r$, $d$ is the local Hilbert space dimension of qudits, and $\text{Wg}_{d^{2}}^{(r)}$ is the Weingarten function with an asymptotic expansion for large $d$ as Zhou and Nahum (2019); Collins and Śniady (2006):
@@
and the weight for the vertical bond (within one two-qudit gate in the original circuit language) is given by the Weingarten function. Besides, mid-circuit measurements can be treated as quenched disorder Weinstein et al. (2022); Liu et al. (2023a), and the diagonal bond with the measurement gives a weight contribution factor $d^{0}$. However, the partition function is ill-defined due to the negative weighted configurations since $\text{Moeb}(\sigma)$, the Moebius number of $\sigma$, can be negative Collins and Śniady (2006). To obtain positive definite weights and properly defined partition function for the statistical model, we integrate out half of the freedoms (all $\tau$ spins) and obtain positive three-body weights of downward triangles as
@@
After tracing out $\tau$, the total weight of a given spin configuration is the product of the weights of all the downward triangles on the space-time lattice of the circuit. In the large $d$ limit, we focus on the most dominant spin configuration of the largest weight, namely, the partition function $Z$ is determined by the weight of the dominant spin configuration. A finite $d$ such as $d=2$ in qubit case introduces interactions between different replicas and a more general entanglement membrane picture is required as discussed in Sec. III.1.
The statistical model is ferromagnetic since the weights of the downward triangles with specific spin configurations as $W^{0}(\sigma,\sigma;\sigma)\approx d^{0}$ and $W^{0}(\sigma^{\prime},\sigma;\sigma)\approx d^{-|(\sigma^{\prime})^{-1})\sigma|}<1$. Therefore, the spin-spin interaction in the effective model is ferromagnetic, and all the spins tend to align in the same direction. However, as discussed below, due to the particular boundary conditions, domain walls may appear, with the unit free energy of $\log\,W^{0}(\sigma^{\prime},\sigma;\sigma)$ for unit length. It is also worth noting that the weight $W^{0}(\sigma,\sigma;\sigma^{\prime})=0$ due to unitary constraint Dong et al. (2021); Li et al. (2023). Consequently, the domain wall in the random unitary circuit is precluded from exhibiting horizontal configurations or upward triangular turning points. However, the presence of mid-circuit measurements and quantum noises Liu et al. (2024a) can relax the unitary constraint, thereby enabling the emergence of such domain walls. These non-unitary factors introduce a new layer of complexity and versatility to the quantum circuit, potentially giving rise to a richer variety of dynamical behaviors.
@@
where $C=\begin{pmatrix}1&2&...&n\\ 2&3&...&1\end{pmatrix}$ and $I=\begin{pmatrix}1&2&...&n\\ 1&2&...&n\end{pmatrix}$ for corresponding qudits, as the cyclic and identity permutations in $S_{n}$ group. With the help of the replica trick Nishimori (2001); Kardar (2007), we can overcome the difficulty of the average outside the logarithmic function
@@
Appendix D Non-Gaussianity of fermionic GHZ states
@@
Appendix E MBL phase transitions in interacting AA model
@@
Appendix F Quasiparticle picture of SSH model

### v3

@@ 5 Discussions
Our settings are different from the conventional TMI settings where all the system qubits are entangled with corresponding reference qubits. Conversely, all qubits in $\bar{E}$ are initialized with pure states in our case. On the one hand, our settings save a great number of qubits ($L+1$ in one-to-one encoding case) and are thus much more friendly for noisy intermediate-scale quantum experiments [139]. On the other hand, the results in our settings rely on the choice of the initial states in some cases. For random quantum circuit evolution, the choice of different product initial states is irrelevant as the difference is smeared out by the first layer of random gates. Instead, the results can be different for different initial states in Hamiltonian quench dynamics. In charge conserved evolution, different choices of the initial states correspond to different charge sectors of different Hilbert space dimensions, which can result in very different behaviors [96]. For different initial states within the same charge sector, the difference is less known. Therefore, it deserves further investigations into the difference in information dynamics from different initial states, particularly in the case of symmetric evolution.
@@
For U(1) symmetric dynamics, the dynamics of mutual information characterized by von Neumann entropy and Rényi entropy could be drastically different. This is because Rényi entropy growth shows subdiffusive behaviors for systems with conservation laws while von Neumann entropy always exhibits ballistic growth [92, 93, 72, 95]. In our work, we focus on von Neumann entropy based mutual information for Hamiltonian dynamics. The dynamical behaviors for Rényi mutual information in these systems as well as the interplay between other symmetries with information dynamics require further investigation.
@@
As previously noted, the framework for information dynamics established within this work is amenable to noisy intermediate-scale quantum [139] technology and is thus well-suited for implementation on near-term quantum computing platforms. Specifically, our settings are more qubit efficient than measuring TMI and are free from Pauli operator average and reversal evolution than measuring OTOC. Given that $A$ and $R$ both being small regions comprising a few qubits, the von Neumann or Rényi entropy for these subsystems can be obtained either via full state tomography or random basis measurement [140, 141, 142]. Apart from experiments on quantum hardware, an interesting theoretical future direction is to analyze the impact of different types of quantum noise on these settings utilizing statistical model mapping. This could lead to the identification of noise-resilient experimental protocols for the implementation of information dynamics. It is also worth noting that random circuit structures utilized in this work are helpful in alleviating coherent quantum noises and demonstrating quantum advantages [143].
@@
This work concentrates on one-dimensional systems, and we expect that higher-dimensional systems have the potential to exhibit a broader range of behaviors in information dynamics. There are also more possible geometry configurations to arrange $E$ and $A$ regions which could lead to a richer understanding of information propagation and scrambling. Another limit is to apply information metrics developed in this work to zero-dimensional systems with all-to-all interactions such as SYK models [144, 145, 146], which yield faster information scrambling and are analytically tractable in the large $N$ limit.
@@
Apart from its implications for non-equilibrium physics, SIC and a deeper understanding of information dynamics can also be helpful and insightful for designing and benchmarking quantum algorithms, particularly in the context of variational quantum algorithms [147, 148]. Promising examples along this direction include understanding the training dynamics and learning capability for variational quantum machine learning algorithms from information perspectives [149, 150, 151]. Besides, it is inspiring to develop better circuit ansatz design [152, 153, 154, 155, 156], parameter initialization [157, 158] and training strategies [159] with enhanced expressive power and alleviated barren plateau issues [160] by leveraging the principles of non-equilibrium dynamics and information spreading.
@@ Acknowledgements
We thank helpful discussions with Zhou-Quan Wan and Shuai Yin. YQC acknowledges the support by NSAF No. U2330401. SXZ acknowledges the support from Innovation Program for Quantum Science and Technology (2024ZD0301700) and the start-up grant at IOP-CAS.
@@ Appendix A Relation between subsystem information capacity and out-of-time correlators
The established relation between mutual information and average OTOC over local Pauli operators [26] can be similarly extended to our case:
@@
The schematic proof based on tensor network representation is shown in Fig. 28. For the tripartite mutual information setting in Ref. [26], the initial state projectors (grey triangles in Fig. 28) are replaced with direct links in pairs since the input state is a fully mixed state for each qubit. In that case, we reproduce the equivalence between Rényi-2 entanglement entropy with the conventional OTOC for local Pauli operators, unlike the introduction of projectors for the pure initial state in our settings.
@@ Appendix C Mapping between random quantum circuits and effective statistical models
The random quantum circuits under investigation are constructed from a sequence of random two-qudit unitary gates, each sampled from the Haar ensemble and arranged in a brickwall layout, as depicted in Fig. 1(a). In this study, we adhere to the convention of defining the unit time $\Delta t$ = 1 for an even-odd layer of two-qudit gates. This convention implies a lightcone speed of $v=2$, which differs from some literature where $\Delta t=1$ is associated with a single even or odd brickwall layer. After $T$ layers of gates, the output density matrix $\rho$ is given by
@@
where $S_{r}$ is the permutation group of degree $r$, $d$ is the local Hilbert space dimension of qudits, and $\text{Wg}_{d^{2}}^{(r)}$ is the Weingarten function with an asymptotic expansion for large $d$ as [75, 78]:
@@
and the weight for the vertical bond (within one two-qudit gate in the original circuit language) is given by the Weingarten function. Besides, mid-circuit measurements can be treated as quenched disorder [64, 65], and the diagonal bond with the measurement gives a weight contribution factor $d^{0}$. However, the partition function is ill-defined due to the negative weighted configurations since $\text{Moeb}(\sigma)$, the Moebius number of $\sigma$, can be negative [78]. To obtain positive definite weights and properly defined partition function for the statistical model, we integrate out half of the freedoms (all $\tau$ spins) and obtain positive three-body weights of downward triangles as
@@
After tracing out $\tau$, the total weight of a given spin configuration is the product of the weights of all the downward triangles on the space-time lattice of the circuit. In the large $d$ limit, we focus on the most dominant spin configuration of the largest weight, namely, the partition function $Z$ is determined by the weight of the dominant spin configuration. A finite $d$ such as $d=2$ in qubit case introduces interactions between different replicas and a more general entanglement membrane picture is required as discussed in Sec. 3.1.
The statistical model is ferromagnetic since the weights of the downward triangles with specific spin configurations as $W^{0}(\sigma,\sigma;\sigma)\approx d^{0}$ and $W^{0}(\sigma^{\prime},\sigma;\sigma)\approx d^{-|(\sigma^{\prime})^{-1})\sigma|}<1$. Therefore, the spin-spin interaction in the effective model is ferromagnetic, and all the spins tend to align in the same direction. However, as discussed below, due to the particular boundary conditions, domain walls may appear, with the unit free energy of $\log\,W^{0}(\sigma^{\prime},\sigma;\sigma)$ for unit length. It is also worth noting that the weight $W^{0}(\sigma,\sigma;\sigma^{\prime})=0$ due to unitary constraint [79, 80]. Consequently, the domain wall in the random unitary circuit is precluded from exhibiting horizontal configurations or upward triangular turning points. However, the presence of mid-circuit measurements and quantum noises [66] can relax the unitary constraint, thereby enabling the emergence of such domain walls. These non-unitary factors introduce a new layer of complexity and versatility to the quantum circuit, potentially giving rise to a richer variety of dynamical behaviors.
@@
where $C=\begin{pmatrix}1&2&...&n\\ 2&3&...&1\end{pmatrix}$ and $I=\begin{pmatrix}1&2&...&n\\ 1&2&...&n\end{pmatrix}$ for corresponding qudits, as the cyclic and identity permutations in $S_{n}$ group. With the help of the replica trick [81, 82], we can overcome the difficulty of the average outside the logarithmic function
@@
Appendix D Experiment protocol advantage for measuring SIC
Beyond its theoretical utility in characterizing quantum dynamics, SIC framework often presents significant practical advantages in terms of experimental implementation compared to other common information scrambling probes like OTOC and TMI. Here, we elaborate on these advantages, with the schematic protocols illustrated in Fig. 29.
Figure 29: Schematic comparison of experimental protocols. (a) A typical measurement protocol for OTOCs, requiring both forward ($U$) and backward ($U^{\dagger}$) time evolution. The figure is taken from Ref. [43] (b) Protocol for measuring SIC (specifically $MI(Q_{a},\{Q_{1},Q_{2}\})$ here), requiring only a single forward time evolution ($U$) followed by subsystem entropy measurements based on random measurement scheme [141].
A primary advantage of SIC measurement lies in its reduced circuit complexity regarding time evolution. As depicted schematically in Fig. 29(a), standard methods for measuring OTOCs typically require executing both the forward time evolution $U$ and the backward time evolution $U^{\dagger}$. Implementing $U^{\dagger}$ accurately can be experimentally demanding, often requiring precise knowledge and decomposition of $U$. The implementation doubles the circuit depth for gate-based evolution, or necessitates fine-tuned pulse reversal in analog systems. This increased complexity increases susceptibility to quantum noise and decoherence.
Measuring TMI, on the other hand, often involves preparing fully mixed initial states where the entire system L is entangled with an equally sized reference system (L). Such thermofield double states require preparing and controlling $2L$ qubits in a highly entangled state.
In stark contrast, the protocol for measuring SIC, particularly the one-to-one encoding scheme predominantly used in this work, is shown in Fig. 29(b), which fundamentally requires only a single application of the forward evolution $U$ and one ancilla qubit. After the system evolves, the task reduces to measuring the necessary subsystem entropies. This elimination of the backward evolution step ($U^{\dagger}$) significantly simplifies the quantum circuit, reduces the overall gate count or evolution time, and consequently minimizes the accumulation of errors.
Finally, the measurement step for SIC involves estimating subsystem von Neumann entropies ($S_{A}$, $S_{R}$, $S_{AR}$). And measuring entanglement entropy requires randomized measurements, where we perform measurements in multiple random basis on copies of the evolved state. For a larger subsystem $|A|>L/2$, we can measure the entanglement entropy of $S_{R\bar{A}}$ instead, to further reduce the measurement overhead, since the whole system is closed.
In conclusion, the requirement of only single forward evolution and the significantly reduced demand on initial entanglement resources contribute to making SIC a potentially more feasible, resource-efficient, and robust probe of quantum information dynamics on current quantum hardware platforms.
Appendix E Non-Gaussianity of fermionic GHZ states
@@
Appendix F MBL phase transitions in interacting AA model
@@
Appendix G Quasiparticle picture of SSH model